# How AI Ideas Affect the Creativity, Diversity, and Evolution of Human Ideas: Evidence From a Large, Dynamic Experiment

### v1

@@ Abstract.
Exposure to large language model output is rapidly increasing. How will seeing AI-generated ideas affect human ideas? We conducted an experiment (800+ participants, 40+ countries) where participants viewed creative ideas that were from ChatGPT or prior experimental participants and then brainstormed their own idea. We varied the number of AI-generated examples (none, low, or high exposure) and if the examples were labeled as “AI” (disclosure). Our dynamic experiment design—ideas from prior participants in an experimental condition are used as stimuli for future participants in the same experimental condition—mimics the interdependent process of cultural creation: creative ideas are built upon prior ideas. Hence, we capture the compounding effects of having LLMs “in the culture loop”. We find that high AI exposure (but not low AI exposure) did not affect the creativity of individual ideas but did increase the average amount and rate of change of collective idea diversity. AI made ideas different, not better. There were no main effects of disclosure. We also found that self-reported creative people were less influenced by knowing an idea was from AI, and that participants were more likely to knowingly adopt AI ideas when the task was difficult. Our findings suggest that introducing AI ideas into society may increase collective diversity but not individual creativity.
creativity, large language models, generative artificial intelligence, chatgpt, co-creation, experiments
copyright: acmcopyrightjournalyear: 2023doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation email; June 03–05, 2024; Woodstock, NYbooktitle: tbdprice: 15.00isbn: 978-1-4503-XXXX-X/18/06ccs: Human-centered computing Human computer interaction (HCI)ccs: Human-centered computing Empirical studies in HCIccs: Human-centered computing Laboratory experimentsccs: Human-centered computing Empirical studies in collaborative and social computingccs: Human-centered computing Computer supported cooperative workccs: Human-centered computing Collaborative content creation
@@ 1. Introduction
If we think of culture as a “loop” where individuals and societies shape each other through exchanges of ideas and practices (Richerson and Boyd, 2008; Boyd and Richerson, 1988), then a fundamental question emerges: What happens when generative AI joins the “culture loop?” This question is not hypothetical. Exposure to LLMs (large language models) is increasing rapidly: When released, ChatGPT was the fastest-growing consumer application in history (Hu, 2023). Moreover, we are likely exposed to even more AI content than we realize: Humans overestimate their ability to distinguish AI from human content (Jakesch et al., 2022). This exposure likely matters: Ideas we see affect the ideas we create (Nijstad and Stroebe, 2006). How, then, will the rapid rise of exposure to LLM-generated ideas affect the creativity, diversity, and evolution of human ideas? And to what extent do AI ideas influence human ideas?
As AI exposure has increased, so have concerns over AI disclosure (Hancock et al., 2020) (whether AI systems should disclose the fact that they are AI). California, for instance, considered passing a law requiring disclosure on behalf of anyone using bots on social media (Williams, 2018). Concerns regarding the disclosure of LLMs are only more likely to grow. In domains ranging from poetry (Köbis and Mossink, 2021) to online social media profiles (Jakesch et al., 2022), LLM output is increasingly indistinguishable from that of humans. We are interested, then, if disclosing ideas as coming from AI moderates the effect of AI exposure.
@@
Our dynamic experiment design—ideas from prior participants are used as stimuli for future participants—mimics the interdependent process of cultural creation: creative ideas are built upon prior ideas. Hence, we capture the compounding effects of having LLMs “in the culture loop”. Our design allows us to observe not just average levels but also temporal dynamics of creativity and diversity in each condition. Taken together, our results provide insights into the role of LLMs in shaping collective thought.
@@
    High AI exposure increases collective diversity but not individual creativity. Passive exposure to AI ideas is important to understand. Due to the interconnected nature of social media, many more people may be exposed to AI content than actively engaged with it. As far as we know, this is the first large-scale experiment to measure how passive exposure to (rather than active engagement with) off-the-shelf LLMs affects collective brainstorming. We find that high AI exposure increases collective idea diversity, but does not affect individual creativity. Our high-powered null finding around creativity can inform public debates over the creative impact of AI ideas on an individual’s creativity. Maybe, our results suggest, there is little impact on individual creativity. However, we found conditions with high levels of AI exposure had more collective idea diversity. That is, ideas in the high AI exposure conditions were more different from each other. Our findings around creativity and diversity suggest the effect of AI exposure may be nuanced: The introduction of AI ideas into human society may yield more diverse but no better human ideas.
@@
    Participants adopted AI ideas more for harder prompts. We find that when AI ideas are disclosed, participants are more likely to adopt the ideas of AI for difficult AUT prompts. This suggests that users will rely on AI ideas not for trivial creative tasks but for difficult ones. But since this finding is based on a small number of prompts, we view this finding as speculative/preliminary and encourage more work on the topic.
@@ 1.1.2. Idea Diversity & AI Adoption
In addition to creativity, we measure how our experimental factors (LLM exposure and LLM disclosure) shape the diversity of ideas that participants produce. This is a complementary measure to creativity. Creativity is often thought of as an individual-level outcome. Diversity is a collective outcome. Put another way, creativity is a property of an idea while diversity is a property of an idea set.  [^4] Hence, it is possible in our framework to have a diverse set of non-creative ideas if each individual idea is not creative (by some metric) but different from one another. We measure two sides of diversity—semantic divergence (which we refer to as idea diversity) and semantic convergence towards AI ideas (which we refer to as AI adoption).
@@ 2. Related Work
Our work bridges two research streams: human-AI co-creation and crowd-sourced creativity. AI ideas are scattered amongst human ideas, whether or not we can tell (Jakesch et al., 2022). This exposure presumably affects the ideas we create (co-creation). And our ideas presumably affect the ideas others create (crowdsourced creativity). After discussing how our study bridges these two streams, we turn to the particular kind of creativity and diversity our experiment captures and what is known about how our two factors (LLM exposure; LLM disclosure) would affect these outcome variables. However, much of the relevant literature gives conflicting predictions, a key motivation for conducting the current study.
2.1. Situating Our Work Between Co-Creation and Crowd Creativity
@@ 2.1.1. Human-AI Co-Creation
As the creative ability of AI advances (Miller, 2019), researchers explored how co-creating with AI affects human creativity. Much of this research explores creative writing with language models, in particular (Gero et al., 2022; Mirowski et al., 2023; Lee et al., 2022; Yang et al., 2022; Yuan et al., 2022; Roemmele, 2021; Di Fede et al., 2022; Hitsuwari et al., 2022; Gero, 2023; Mizrahi et al., 2020; Gero and Chilton, 2019). While most prior work in this domain involves users actively engaging with custom systems, our study is concerned with passive exposure to outputs from off-the-shelf models. (By ‘passive exposure’ we mean that (1) users are shown LLM outputs but did not have an active role in the creation of these outputs and that (2) users were given no instructions to actively engage with these outputs; they were merely shown the LLM outputs. ‘Passive exposure’ approximates how users often encounter LLM outputs in the real world.) Human-AI co-creation shows that the relationship between AI ideas and their effect on human creativity is nuanced, with task-level and attitudinal factors playing a role. Roemmele (2021) found seeing AI examples influenced outcomes for hard, but not easy, prompts. Gero et al. (2022) found the quality of LLM outputs did not correlate with perceived usefulness. This is consistent with other research showing large variance in the perceived usefulness of outputs from co-creation systems (Calderwood et al., 2020), suggesting human attitudes partially determine the utility of AI creativity aids. We extend this predominantly qualitative work with a large-scale quantitative study.
Several studies (Wieland et al., 2022; Hwang and Won, 2021; Salikutluk et al., 2023) explored the most relevant co-creation task to our study: brainstorming with chatbots. Hwang and Won (2021) developed a study design where participants generated ideas with a partner who was either a human confederate or a scripted chatbot. Participants produced more creative ideas when they believed their partner was a chatbot. Hwang and Won (2021) explain this using social facilitation theory—the perceived novelty of collaborating with an AI system enhanced performance. This core finding was replicated in Wieland et al. (2022). We build on this work by viewing the effect of AI disclosure through social facilitation. But our study has two key differences from Hwang and Won (2021). First, we consider exposure to real AI ideas rather than a scripted chatbot. Second, we focus on passive exposure and not direct engagement; perhaps the novelty of AI is lower when users are passively exposed to its ideas vs. directly engaging with it.
@@ 2.1.3. Our Contributions
Our study incorporates elements of human-AI co-creation and crowd creativity, complementing both fields. We note that co-creation studies often confound the effect of exposure with the effect of disclosure: If one is creating with an AI system, it is impossible to separate the content of an AI system from the knowledge that the content is from an AI system. Our factorial design lets us estimate the marginal effect of exposure and disclosure separately. Co-creation studies typically employ a small number of specialized participants actively engaged with a system. From the perspective of validating a system, this is reasonable. But we are interested in the effects of (1) passive exposure on (2) a general public. For this reason, we adopt a large-scale experimental design—similar to crowd-sourced creativity studies—that lets us estimate effects on the general public rather than specialized users. A key benefit of our large sample size is that we can precisely estimate how participant attitudes affect human-AI outcomes. This is important because, as Gero et al. (2022, pg. 1016) write: “[P]articipant attitudes are a major unknown factor when studying human-AI collaboration.” Our experiment design also lets us understand the effect of AI over time since responses feed forward, allowing us to observe differences in rates of change between conditions.
@@ 2.2.1. Effects on Creativity
Intuitively, the effect of exposure to ChatGPT ideas will depend on how creative ChatGPT answers are relative to human ideas. In preliminary testing, we found that the answers to the AUT generated by our prompt were scored as roughly 12 percentage points more creative than the ideas generated by humans (see Appendix E) via the Organisciak et al. (2022) classifier. LLM generations may be increasing in creativity: while GPT-3 (an earlier model than ChatGPT-3.5) scored lower in AUT creativity than humans on the AUT (Stevenson et al., 2022), GPT-4 (a more recent model than ChatGPT-3.5) scored among the top percentile of humans on a similar verbal creativity task (of Minnestota, [n. d.]).
Even if language models can generate creative ideas, it is unclear from prior work if exposure to these ideas can increase human creativity. On one hand, the associative model of brainstorming suggests that exposure to others’ ideas can stimulate idea generation by activating a non-accessible concept of a participant’s memory (Nijstad and Stroebe, 2006; Brown and Paulus, 2002; Paulus and Brown, 2007). For example, ChatGPT may come up with a use for a bottle that you never associated with bottles. This can then inspire you to come up with creative uses along this line. In this way, ChatGPT can stimulate creativity. On the other hand, there is also evidence that seeing the ideas of others inhibits a participant’s idea generation if “one is exposed to an idea that has few connections to other ideas in an individual’s semantic network” (Paulus and Brown, 2007, pg. 10). Indeed, this appeared to be the case in Yang et al. (2022). There is a possibility that AI ideas are creative but so divorced from how humans generate ideas that seeing these ideas actually has an inhibiting effect. Separate from prior academic work, there are public debates about the impact of LLMs (such as ChatGPT) on creativity (e.g., (News, 2023; Review, 2023; Jared Henderson, 2022; Krish Naik, 2023; Tubefilter, 2023; Eapen et al., 2023; Wilcot, 2023)). Many of these debates assume ChatGPT will have some impact on an individual’s creativity—either good or bad. Our work contributes empirical results to this broader public conversation.
@@ 2.2.2. Effects on Diversity
Prior work in AI co-creation finds that actively collaborating with AI can lead to more diverse or unexpected outputs (Yang et al., 2022; Osone et al., 2021; Lee et al., 2022; Branch et al., 2021; Gero and Chilton, 2019). But note that these studies are testing active engagement with intentionally constructed systems. This is different from the passive, incidental exposure to AI ideas that now occurs in everyday life. Writers call ChatGPT ‘a blurry JPEG of the internet’ (Chiang, 2023) and discuss its ‘incredible blandness’ (Mangalaseril, 2023); researchers call it a ‘stochastic parrot’ (Bender et al., 2021). It is not clear, then, if ideas from off-the-shelf LLMs—precisely the kind we are inundated with—also increase the diversity of human ideas.

### v2

@@ Abstract.
Exposure to large language model output is rapidly increasing. How will seeing AI-generated ideas affect human ideas? We conducted a dynamic experiment (800+ participants, 40+ countries) where participants viewed creative ideas that were from ChatGPT or prior experimental participants, and then brainstormed their own idea. We varied the number of AI-generated examples (none, low, or high exposure) and if the examples were labeled as “AI” (disclosure). We find that high AI exposure (but not low AI exposure) did not affect the creativity of individual ideas but did increase the average amount and rate of change of collective idea diversity. AI made ideas different, not better. There were no main effects of disclosure. We also found that self-reported creative people were less influenced by knowing an idea was from AI and that participants may knowingly adopt AI ideas when the task is difficult. Our findings suggest that introducing AI ideas may increase collective diversity but not individual creativity.
artificial intelligence, creativity, large language models, cultural evolution
copyright: acmlicensedjournalyear: 2024doi: XXXXXXX.XXXXXXXconference: Preprint; Jul 03, 2024; Woodstock, NYisbn: 978-1-4503-XXXX-X/18/06ccs: Human-centered computingccs: Human-centered computing Human computer interaction (HCI)ccs: Human-centered computing Empirical studies in HCIccs: Human-centered computing Empirical studies in collaborative and social computingccs: Computing methodologies Artificial intelligenceccs: Computing methodologies Natural language processingccs: Human-centered computing Collaborative interaction
@@ 1. Introduction
If we think of culture as a “loop” where individuals and societies shape each other through exchanges of ideas and practices (Richerson and Boyd, 2008; Boyd and Richerson, 1988), then a question emerges: What happens when generative AI enters the “culture loop?” Exposure to LLMs (large language models) is increasing rapidly: When released, ChatGPT was the fastest-growing consumer application in history (Hu, 2023). Moreover, we are likely exposed to even more AI content than we realize: Humans overestimate their ability to distinguish AI from human content (Jakesch et al., 2022). This exposure likely matters: Ideas we see affect the ideas we create (Nijstad and Stroebe, 2006). How will the rapid rise of exposure to LLM-generated ideas affect the creativity, diversity, and evolution of human ideas? And to what extent do AI ideas influence human ideas?
The scale of ‘passive exposure’ to AI ideas is high, and different from prior human-AI interactions. By ‘passive exposure’, we refer to cases when (A) users see LLM outputs but do not have an active role in the creation of these outputs and (B) users are given no instructions to actively engage with these outputs. ‘Passive exposure’ approximates how users often encounter LLM outputs in the real world. For example, OpenAI users generate 100 billion words per day (Griffin, 2024). It is likely that the number of people who are merely seeing (i.e., passively exposed) AI output is significantly larger than the number of people who are creating (i.e., actively engaging) with these systems. Arguably, many future human-AI teams will exhibit this relationship. Yet in existing studies of human-AI creativity, participants are often actively interacting with an AI system (Yang et al., 2022; Osone et al., 2021; Lee et al., 2022; Branch et al., 2021; Gero and Chilton, 2019; Padmakumar and He, 2024). It is crucial, then, to understand how passive AI exposure shapes human ideas.
As AI exposure has increased, so have concerns over AI disclosure (Hancock et al., 2020) (whether providers should disclose when they use AI systems). California, for instance, considered passing a law requiring disclosure on behalf of anyone using bots on social media (Williams, 2018). Concerns regarding the disclosure of LLMs are only more likely to grow. In domains ranging from poetry (Köbis and Mossink, 2021) to online social media profiles (Jakesch et al., 2022), LLM output is increasingly indistinguishable from that of humans. We are interested, then, if disclosing ideas as coming from AI moderates the effect of AI exposure.
@@
Our dynamic experiment design—ideas from prior participants are used as stimuli for future participants—speaks to the interdependent process of cultural creation: creative ideas are built upon prior ideas. Hence, we capture the compounding effects of having LLMs “in the culture loop”. It is also intended to mimic possible futures for human-AI teams. Our design allows us to observe not just average levels but also temporal dynamics of creativity and diversity in each condition. Taken together, our results provide insights into the role of LLMs in shaping collective thought.
@@
    High AI exposure increases collective diversity but not individual creativity. We find that high AI exposure increases collective idea diversity, but does not affect individual creativity. Our high-powered null finding around creativity can inform public debates over the creative impact of seeing AI ideas ‘in the wild’. However, we found conditions with high levels of AI exposure had more collective idea diversity. That is, ideas in the high AI exposure conditions were more different from each other. Our findings around creativity and diversity suggest the effect of AI exposure may be nuanced: The introduction of AI ideas into human society may yield more diverse but no better human ideas.
@@
    Participants may adopt AI ideas for harder prompts. We find that when AI ideas are disclosed, participants are more likely to adopt the ideas of AI for difficult AUT prompts. This suggests that users will rely on AI ideas not for trivial creative tasks but for difficult ones. But since this finding is based on a small number of prompts, we view this finding as speculative and encourage more work on the topic.
@@ 1.1.2. Idea Diversity & AI Adoption
In addition to creativity, we measure how our experimental factors (LLM exposure and LLM disclosure) shape the diversity of ideas that participants produce. This is a complementary measure to creativity. Creativity is often thought of as an individual-level outcome. Diversity is a collective outcome. Put another way, creativity is a property of an idea while diversity is a property of an idea set. We measure two sides of diversity—semantic divergence (which we refer to as idea diversity) and semantic convergence towards AI ideas (which we refer to as AI adoption).
@@ 2. Related Work
Our work bridges three research streams: human-AI co-creation, crowd-sourced creativity, and complex systems. AI ideas are scattered amongst human ideas, whether or not we can tell (Jakesch et al., 2022). This exposure presumably affects the ideas we create (co-creation). And our ideas presumably affect the ideas others create (crowdsourced creativity). While real-world culture is dynamic and evolving, most experiments are not set up to capture evolution (collective dynamics). By employing a large-N sample size and ‘many-worlds’ setup, we model the complex dynamics of AI influence. After discussing how our study bridges these streams, we turn to the particular kind of creativity and diversity our experiment captures and what is known about how our two factors (LLM exposure; LLM disclosure) would affect these outcome variables. However, much of the relevant literature gives conflicting predictions, a key motivation for conducting the current study.
2.1. Situating Our Work Between Co-Creation, Crowd Creativity, and Collective Dynamics
@@ 2.1.1. Human-AI Co-Creation
As the creative ability of AI advances (Miller, 2019), researchers explored how co-creating with AI affects human creativity. Much of this research explores creative writing with language models, in particular (Gero et al., 2022; Mirowski et al., 2023; Lee et al., 2022; Yang et al., 2022; Yuan et al., 2022; Roemmele, 2021; Di Fede et al., 2022; Hitsuwari et al., 2022; Gero, 2023; Mizrahi et al., 2020; Gero and Chilton, 2019; Padmakumar and He, 2024). While most prior work in this domain involves users actively engaging with custom systems, our study is concerned with passive exposure to outputs from off-the-shelf models. (By ‘passive exposure’ we mean that (1) users are shown LLM outputs but did not have an active role in the creation of these outputs and that (2) users were given no instructions to actively engage with these outputs; they were merely shown the LLM outputs. ‘Passive exposure’ approximates how users often encounter LLM outputs in the real world.) Human-AI co-creation shows that the relationship between AI ideas and their effect on human creativity is nuanced, with task-level and attitudinal factors playing a role. Roemmele (2021) found seeing AI examples influenced outcomes for hard, but not easy, prompts. Gero et al. (2022) found the quality of LLM outputs did not correlate with perceived usefulness. This is consistent with other research showing large variance in the perceived usefulness of outputs from co-creation systems (Calderwood et al., 2020), suggesting human attitudes partially determine the utility of AI creativity aids. We extend this predominantly qualitative work with a large-scale quantitative study.
Most similar to our work, several studies have explored how (post-ChatGPT) generative artificial intelligence affects creativity and diversity. Several studies found that ideating with generative AI can decrease diversity (Anderson et al., 2024; Doshi and Hauser, 2024; Padmakumar and He, 2024). Some studies suggest generative AI increases individual creativity (Doshi and Hauser, 2024; Dell’Acqua et al., 2023) while others (Anderson et al., 2024) find no effect. Our study offers several additions to this literature. First, because of its dynamic design, we test the long-run effects of AI, where ideas feed forward to future participants. Second, our study is concerned with “passive exposure”: Participants are not told ideas are from AI, and are not instructed to engage with these ideas. By systematically ablating whether AI ideas are disclosed as such, we can explore if the effect of AI ideas depends on knowledge of where the idea is from. Third, we employ a large sample size—which is useful since it provides power for precise estimates of effects and the ability to capture heterogeneity. Moreover, our large sample is comprised of creative professionals and technology-oriented users, two groups most relevant to the phenomenon.
@@ 2.1.2. Crowdsourced Creativity
2.1.3. Collective Dynamics & Many-Worlds Experiments
Prior work in complex systems and computational sociology highlights the importance of studying collectives to understand social dynamics (Salganik and Watts, 2009). Meanwhile, traditional experiments focus on individuals. Identifying the effect of AI ideas on the diversity and evolution human ideas similarly requires an examination of complex systems as opposed to individuals in isolation. Hence, our experimental design draws on the “many-world” paradigm (e.g., (Salganik et al., 2006)):We create multiple, parallel realizations of worlds with and without AI ideas, each evolving independently under controlled conditions. By employing a large-N sample size and many different parallel worlds, we can better understand the collective effects of AI ideas on human ideas. Beyond simple averages, we can also model how AI ideas affect the evolution of human ideas.
@@ 2.1.4. Our Contributions
Our study incorporates elements of human-AI co-creation, crowd creativity, and collective dynamics. We note that co-creation studies often confound the effect of exposure with the effect of disclosure: If one is creating with an AI system, it is impossible to separate the content of an AI system from the knowledge that the content is from an AI system. Our factorial design lets us estimate the marginal effect of exposure and disclosure separately. Co-creation studies typically employ a small number of specialized participants actively engaged with a system. From the perspective of validating a system, this is reasonable. But we are interested in the effects of (1) passive exposure on (2) a general public. For this reason, we adopt a large-scale experimental design—similar to crowd-sourced creativity studies—that lets us estimate effects on the general public rather than specialized users. A key benefit of our large sample size is that we can precisely estimate how participant attitudes affect human-AI outcomes. This is important because, as Gero et al. (2022, pg. 1016) write: “[P]articipant attitudes are a major unknown factor when studying human-AI collaboration.” Drawing on the “many-worlds” paradigm, our experiment design also lets us understand the effect of AI over time since responses feed forward, allowing us to observe differences in rates of change between conditions.
@@ 2.2.1. Effects on Creativity
Intuitively, the effect of exposure to ChatGPT ideas will depend on how creative ChatGPT answers are relative to human ideas. In preliminary testing, we found that the answers to the AUT generated by our prompt were scored as more creative than the ideas generated by humans (see Appendix E) via the Organisciak et al. (2022) classifier. LLM generations may be increasing in creativity: while GPT-3 (an earlier model than ChatGPT-3.5) scored lower in AUT creativity than humans on the AUT (Stevenson et al., 2022), GPT-4 (a more recent model than ChatGPT-3.5) scored among the top percentile of humans on a similar verbal creativity task (Guzik et al., 2023), as measured by human judges.
Even if language models can generate creative ideas, it is unclear from prior work if mere exposure to these ideas can increase human creativity. On one hand, the associative model of brainstorming suggests that exposure to others’ ideas can stimulate idea generation by activating a non-accessible concept of a participant’s memory (Nijstad and Stroebe, 2006; Brown and Paulus, 2002; Paulus and Brown, 2007). For example, ChatGPT may come up with a use for a bottle that you never associated with bottles. This can then inspire you to come up with creative uses along this line. In this way, ChatGPT can stimulate creativity. On the other hand, there is also evidence that seeing the ideas of others inhibits a participant’s idea generation if “one is exposed to an idea that has few connections to other ideas in an individual’s semantic network” (Paulus and Brown, 2007, pg. 10). Indeed, this appeared to be the case in Yang et al. (2022). There is a possibility that AI ideas are creative but so divorced from how humans generate ideas that seeing these ideas actually has an inhibiting effect. Separate from prior academic work, there are public debates about the impact of LLMs (such as ChatGPT) on creativity (e.g., (News, 2023; Review, 2023; Jared Henderson, 2022; Krish Naik, 2023; Tubefilter, 2023; Eapen et al., 2023; Wilcot, 2023)). Many of these debates assume ChatGPT will have some impact on an individual’s creativity—either good or bad. Our work contributes empirical results to this broader public conversation.
@@ 2.2.2. Effects on Diversity
Prior work in AI co-creation finds mixed effects. Collaborating with AI can lead to more diverse (Yang et al., 2022; Osone et al., 2021; Lee et al., 2022; Branch et al., 2021; Gero and Chilton, 2019) or less diverse outputs (Padmakumar and He, 2024; Doshi and Hauser, 2024; Dell’Acqua et al., 2023). But note that these studies are testing active engagement, and most test active engagement with intentionally constructed systems. This is different from the passive, incidental exposure to AI ideas that now occur in everyday life. Writers call ChatGPT ‘a blurry JPEG of the internet’ (Chiang, 2023) and discuss its ‘incredible blandness’ (Mangalaseril, 2023); researchers call it a ‘stochastic parrot’ (Bender et al., 2021). It is not clear, then, how passive exposure to ideas from off-the-shelf LLMs—precisely the kind we are inundated with—would affect the diversity of human ideas.